### v2

@@ 3.6.2. Evaluation Models
Our goal is to evaluate the foundational capabilities of SOTA LLMs on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, to better understand their strengths and limitations in financial tagging. To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Table 13. These include: (1) one general-purpose closed-source LLM: GPT-4o model (Hurst et al., 2024); (2) eleven general-purpose open-source LLMs including DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025), Qwen3-32B (Team, 2025), Qwen3-14B (Team, 2025), Qwen3-1.7B (Team, 2025), Qwen3-0.6B (Team, 2025), Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), and gemma-2-27b-it (Team, 2024) models; and (3) one domain-specific financial LLM: Fino1-8B (Qian et al., 2025) model. In addition, we compare these LLMs with strong PLMs, including BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022).
3.6.3. Evaluation setting
We use the LM Evaluation Harness (Gao et al., 2024) to build customized benchmark suites. Proprietary models are accessed via APIs, while open-source models are evaluated locally on a 4×H100 GPU cluster (80GB each). We standardize the input length to 2,048 tokens for FinNI and 4,096 for FinCL, with a generation limit of 1,024 tokens to support reasoning-intensive outputs. For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix I). All data follow the FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) formats. For retrieval in FinCL, we use ElasticSearch [^8] https://www.elastic.co/ with text-embedding-3-small as the embedding model. For evaluation, we spent a total of 500 GPU hours and about \$1,500 for GPT (from OpenAI) and DeepSeek (from TogetherAI) API.
4. Experiment and Result
4.1. Overall Results
@@
Table 5. Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@
From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3, GPT-4o, and Llama-4-Scout-17B-16E-Instruct achieve the top 3 macro-F1 scores (0.0659, 0.0569, and 0.0400), outperforming all fine-tuned PLMs. This highlights the strong generalization of large LLMs and the effectiveness of our task design. Qwen3-32B also achieves a solid macro-F1 (0.0324), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings. From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1162, ranking third overall despite no fine-tuning. GPT-4o also performs competitively with a score of 0.0863, outperforming most open-source and domain-specific models.
4.2. Subtask Results
4.2.1. FinNI subtask
Table 6 presents the results of different models on the FinNI subtask. Overall, larger models demonstrate stronger capabilities in identifying numerical entities and producing structured outputs, even without financial domain adaptation. Among them, DeepSeek-V3 achieves the best performance with both the highest recall (0.8430) and F1 (0.6932), while also maintaining competitive precision. In comparison, GPT-4o, the closed-source baseline, performs relatively well with balanced scores, securing the second-best F1 (0.5397). Interestingly, Qwen3-32B and Qwen3-14B yield very high precision (0.6991 and 0.6912, respectively), but their recall is extremely low, leading to weak F1 scores, which highlights that precision alone is insufficient in this task.
Table 6. Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.1. FinNI subtask
Smaller open-source models such as Qwen3-1.7B and Qwen3-0.6B show severe limitations, with F1 below 0.1, underscoring the difficulty of the FinNI subtask in zero-shot settings. In contrast, models like gemma-2-27B-it achieve relatively balanced results, ranking third in both recall and F1. Finally, the domain-specific Fino1-8B, despite being fine-tuned on financial reasoning QA, underperforms compared to general-purpose LLMs. This suggests that domain pretraining on financial text alone provides limited benefit unless the training task aligns closely with FinNI’s requirements.
4.2.2. FinCL subtask
Table 7. Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.2. FinCL subtask
Table 7 reports the accuracy of different models on the FinCL subtask, which is performed on candidate sets obtained from the best top-200 retrieval results, as shown in Table 16 (Appendix J). Although the retrieval stage provides a foundation, the low Acc@k values, especially for table-based entities, directly constrain the FinCL performance.
Within this reranking formulation, DeepSeek-V3 achieves the best accuracy at 0.1889, closely followed by GPT-4o at 0.1829 and Llama-4-Scout-17B-16E-Instruct at 0.1649. Other large open-source models perform notably worse, with Llama-3.3-70B at 0.1318. The medium-scale models, such as Qwen3-32B (0.1277), DeepSeek-R1-Distill-Qwen-32B (0.1141), gemma-2-27B-it (0.1099), and Qwen3-14B (0.1144), show similar levels. While smaller models, including Llama-3.2-3B and Qwen3-0.6B, perform near random with 0.0415 and 0.0414. The domain-specific Fino1-8B also lags at 0.0807, suggesting that domain pretraining without explicit task alignment offers little benefit. As a result, overall accuracy remains low, reflecting both the inherent difficulty of handling complex taxonomies and subtle financial semantics, and the cascading effect of retrieval errors that further lowers FinCL performance.
4.3. Ablation analysis
We further compare our benchmark against extreme multi-class classification settings. The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category (closed-source, open-source, and financial LLMs). The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
Table 8. Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
@@ 4.3. Ablation analysis
Table 8 shows that extreme classification is not a valid way to evaluate LLMs on XBRL tagging. This single-step formulation requires presenting the full space of thousands of US-GAAP concepts and directly assigning tags to raw text, without first identifying which spans are true numerical entities. In practice, this is infeasible: the model cannot realistically process the entire label space and is forced to rely only on its internal knowledge, which leads to meaningless zero-shot predictions. As the results indicate, all tested models, including GPT-4o and DeepSeek-V3, collapse to zero precision, recall, and F1 under this setting.
In contrast, our 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark reformulates the task into a two-stage process that aligns with how tagging is actually performed. The FinNI first detects and types coarse-grained numerical entities from text and tables, ensuring that only valid values enter the linking step. The FinCL then narrows the taxonomy search space by retrieving a candidate set of relevant tags and requiring the LLM to select the best match using contextual reasoning. With this design, models achieve non-trivial performance: for example, DeepSeek-V3 reaches an F1 of 0.1356 and GPT-4o 0.1026. These results demonstrate that FinTagging provides a more faithful evaluation of LLMs’ zero-shot tagging capabilities, surfacing meaningful signals where extreme classification fails.
4.4. Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
Table 9 breaks down model accuracy across three evaluation settings: full pipeline, FinCL, and reranking-only in FinCL with guaranteed gold concepts in the top 200 candidates. The results reveal clear evidence of cascading errors. End-to-end pipeline accuracy is consistently the lowest, since errors from entity extraction (FinNI) and retrieval both propagate into concept linking. When gold entities are provided, accuracy improves substantially (e.g., GPT-4o from 0.1052 to 0.1829; DeepSeek-V3 from 0.1074 to 0.1889), confirming that entity identification is a key bottleneck in this context. With gold candidates guaranteed in the pool (rerank-only), performance further improves (e.g., GPT-4o 0.2023; DeepSeek-V3 0.2144), showing that narrowing the label space alleviates but does not eliminate difficulty, as semantic ambiguity among fine-grained US-GAAP concepts remains challenging.
Table 9. Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
@@
4.5. Error cases
Table 10. Error cases from DeepSeek-V3 rerank.
Case 1: \$31m
Gold Concept Rank 18 / 200 (not in Top-5)
Case 2: \$121m
Gold Concept Rank 2 / 200 (in Top-5)
5. Conclusion
@@
This paper presents 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports. The task is divided into two subtasks, financial numeric identification (FinNI) and concept linking (FinCL), to enable fine-grained evaluation of both information extraction and taxonomy alignment. Results show that while LLMs generalize well to long-tail entities and perform competitively in zero-shot settings, they struggle with accurate alignment to GAAP concepts. This reveals limitations in structure-aware reasoning and highlights the need for better semantic understanding. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: offers a foundation for advancing research in XBRL tagging and regulatory reporting. The limitations of our work and directions for future research are discussed in Appendix A.
Appendix A Limitations
@@
In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated the latest models, such as GPT-5, gpt-oss, and Qwen3-235B. We are currently conducting these experiments and will report the results in future work. Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based financial tagging agents through targeted fine-tuning to further enhance performance.
Appendix B Significance analysis

### v3

@@ 3.6.2 Evaluation Models
Our goal is to assess the foundational capabilities of SOTA LLMs on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and understand their strengths and limitations in XBRL tagging. We evaluate models across three categories using our unified framework (Table 11). (1) Closed-source general LLM: GPT-4o hurst2024gpt. (2) Open-source general LLMs: DeepSeek-V3 liu2024deepseek, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability, Qwen3-32B / 14B / 1.7B / 0.6B qwen3technicalreport, Llama-4-Scout-17B-16E-Instruct meta2025llama, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct grattafiori2024llama, and gemma-2-27b-it gemma_2024. (3) Domain-specific financial LLM: Fino1-8B qian2025fino1. We also compare against strong PLM baselines: BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer.
3.6.3 Evaluation setting
[…]
4 Experiment and Result
4.1 Overall Results
@@
Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@
From a macro perspective, which emphasizes balanced performance across frequent and rare tags, large general-purpose LLMs clearly dominate. DeepSeek-V3, GPT-4o, and Llama-4-Scout achieve the strongest macro-F1 scores, surpassing all fine-tuned PLMs and indicating better generalization to long-tail concepts. Qwen3-32B also performs competitively, suggesting that model architecture and pretraining can compensate for a smaller scale. From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage (Table 5).
4.2 Subtask Results
4.2.1 FinNI subtask
Table 6 summarizes results on FinNI. Overall, larger LLMs show markedly stronger ability to identify numerical entities and produce structured outputs, even without financial adaptation. DeepSeek-V3 achieves the best overall performance, while GPT-4o performs competitively and ranks among the top models.
Table 6: Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.1 FinNI subtask
A notable pattern is that some models (e.g., Qwen3 variants) obtain very high precision but extremely low recall, indicating that they extract only a small fraction of target facts. This confirms that precision alone is insufficient for FinNI, where missing facts are often more damaging than producing a few spurious ones.
Smaller open-source models struggle severely, reflecting the difficulty of this zero-shot extraction setting. In contrast, mid-size models such as gemma-2-27B achieve more balanced behavior. Interestingly, the financial LLM Fino1-8B still lags behind general-purpose models, suggesting that domain text pretraining alone does not help unless the training task aligns closely with FinNI.
4.2.2 FinCL subtask
Table 7 reports FinCL performance under the reranking setup, using candidates retrieved from the best top-200 configuration (Table 14). Although retrieval provides essential coverage, low Acc@k values, especially for table entities, substantially cap downstream accuracy.
Table 7: Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.2 FinCL subtask
Within this setting, large LLMs such as DeepSeek-V3, GPT-4o, and Llama-4-Scout perform best, but overall accuracy remains low, reflecting the difficulty of aligning evidence with fine-grained taxonomy concepts. Most other models cluster well below this tier, and smaller models approach random behavior. The financial model Fino1-8B also lags behind general LLMs, suggesting that domain pretraining alone is insufficient without task-aligned supervision.
Taken together, FinCL exposes two compounding challenges: taxonomy disambiguation requires reasoning over subtle semantic cues, and retrieval errors propagate directly into reranking, creating a ceiling effect on end-to-end alignment.
4.3 Ablation analysis
We further compare our benchmark against extreme multi-class classification settings. The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category. The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
Table 8: Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
@@ 4.3 Ablation analysis
As shown in Table 8, treating XBRL tagging as single-step extreme classification fails in practice. The model must choose directly from thousands of taxonomy concepts without first identifying valid numerical entities, forcing it to rely on internal priors rather than grounded evidence. All models collapse to zero, illustrating that this formulation is not a meaningful way to evaluate LLMs on XBRL tagging. In contrast, our two-stage framework mirrors how tagging is actually performed. FinNI first detects and types candidate numerical facts, and FinCL then retrieves a small set of plausible tags and requires the model to select among them using context. Under this formulation, models achieve non-trivial performance, revealing real differences in capability that are completely hidden under extreme classification.
4.4 Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
Table 9 compares accuracy across three settings: the full pipeline, FinCL with gold entities, and rerank-only with guaranteed gold candidates. The pattern reveals clear cascading errors. End-to-end performance is consistently the lowest because mistakes from extraction and retrieval accumulate in the linking stage.
When gold entities are supplied, accuracy improves substantially, showing that entity identification is a major bottleneck. Performance increases again in the rerank-only setting, indicating that narrowing the label space reduces difficulty. However, accuracy still remains far from solved, suggesting that fine-grained US-GAAP disambiguation is intrinsically challenging even when retrieval is perfect. In addition, we provide representative error cases in Appendix I, illustrating how these challenges arise in practice.
Table 9: Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
@@
5 Conclusion
@@
This paper introduces 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating LLMs on XBRL tagging of real financial reports. The task is structured into two subtasks, Financial Numerical Identification and Financial Concept Linking, to separately assess extraction and concept alignment. Our results show that LLMs perform reasonably well in zero-shot settings and generalize to long-tail entities, but still struggle to align facts with precise US-GAAP concepts. These findings point to gaps in structure-aware and semantic reasoning. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a foundation for future work on reliable XBRL tagging and regulatory reporting.
Limitations
@@
In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated the latest models, such as GPT-5.2, gpt-oss, and Qwen3-235B. We are currently conducting these experiments and will report the results in future work. Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based XBRL tagging agents through targeted fine-tuning to enhance performance further.
Potential Risks
Potential Positive Societal Impacts.
0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first structure-aware benchmark for evaluating large language models on realistic XBRL tagging. By reframing tagging as a transparent, two-stage Extract–Align process, this benchmark:
  • Enables more rigorous and reproducible evaluation of model performance on financial fact extraction and taxonomy alignment.
  • Supports research on safer automation of financial workflows by exposing failure modes that arise in long-tailed concept spaces and complex tabular contexts.
  • Lowers the barrier for auditing, interpreting, and improving financial AI systems by releasing open datasets, evaluation protocols, and diagnostic analyses.
By emphasizing open evaluation rather than proprietary pipelines, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: contributes to more transparent and accountable financial AI research.
Potential Negative Societal Impacts.
Despite these benefits, several risks remain:
  • Misuse and Overreliance. High benchmark performance may be mistakenly interpreted as readiness for real-world XBRL tagging, regulatory reporting, or financial decision-making, even though the benchmark does not replace human review or auditing.
  • Bias and Coverage Limitations. The dataset is drawn from U.S. SEC filings and US-GAAP concepts. As a result, reporting practices, sector distributions, and taxonomy frequency patterns may introduce biases and may not generalize to other jurisdictions or accounting standards.
  • Benchmark Narrowness. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluates extraction and alignment only. Downstream tasks such as aggregation, reconciliation, audit validation, and materiality assessment remain outside the benchmark and may still fail in deployed systems.
Ongoing iteration, external feedback, and careful documentation are necessary to mitigate these concerns.
Ethical Considerations and Usage Disclaimer
The authors assume responsibility for the design and dissemination of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. All data used in the benchmark are derived from publicly available SEC filings and contain no intentionally included personal or sensitive information. Dataset construction followed community ethical guidelines for research involving public financial disclosures.
Unless otherwise specified, the datasets and accompanying evaluation resources are released for research and educational purposes. They should not be used as a substitute for professional accounting judgment, financial reporting systems, or regulatory compliance procedures.
The benchmark is designed to evaluate model behavior under controlled conditions. Performance on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: does not imply suitability for production deployment, financial auditing, or automated investment or compliance tasks. Any real-world use must include human oversight, documentation of limitations, and appropriate legal and regulatory review.
We used AI-assisted writing tools to help refine the wording and improve the readability of the manuscript. All scientific claims, dataset design choices, and experimental analyses were developed and verified by the authors.
While reasonable care has been taken to ensure accuracy, the authors provide no warranty regarding completeness or suitability for any particular purpose and disclaim liability for any damages arising from use of the benchmark or associated materials.
Licensing and Terms of Use
We release all benchmark data, annotations, and accompanying resources under the CC-BY 4.0 license, which permits redistribution and adaptation with attribution. All source documents originate from publicly available SEC filings, and we strictly comply with the original terms governing their public distribution. The US-GAAP taxonomy is used in accordance with its publicly documented terms. Our released benchmark does not contain proprietary material beyond what is already publicly accessible, and any derivative use must continue to acknowledge both the original providers and our benchmark.
Appendix A Significance analysis